LHC 2011 - latest news

Week 44: Machine coordinators: G. Arduini, R. Assmann, MD Coordinators: R. Assmann, F. Zimmermann

Goals of the week: MDs, start setting-up with ions




Link to report from morning meeting

Monday 7th November

  • Night: loss maps. ALFA roman pot could not be moved in position.
  • Transfer line steering optimization
  • 06:30 Start of technical stop.

08:30 meeting (week summary)

Heavy ion commissioning

Sunday 6th November

  • 02:00 Collisions in point 1/2/5. Quiet beams for 1 hour
  • 05:30-06:30 Loss maps at injection
  • 06:30-09:30 Collimator (TCT) control problem in point 1. Access
  • Still problem with collimator in point 1 requiring another access
  • 16:30-19:00 TCT setting-up in point 2 separated and in collision
  • Night: loss maps. ALFA roman pot could not be moved in position.

09:00 meeting

Saturday 5th November

  • 02:00 Start MD on quench margin at injection
  • 05:30 Start BI MD
  • 09:30 Start of BIMD
  • 14:30 RQT trip - end of MD
  • Change polarity of Solenoid/Dipole for ALICE. Switch OFF LHCb dipole.
  • 17:00 Start ion commissioning
  • 21:30 Both beams circulating and captured

09:00 meeting

IR2 Aperture measurements summary (M. Giovannozzi)

Friday 4th November

  • No beam form SPS due to beam dump kicker problem. Investigation on aperture issues L2 continuing (Survey)
  • 19: 30 Beam back. Start preparation for MD on tight collimator settings

08:30 meeting

Thursday 3rd November

  • 09:30 In shadow of SPS recovery: Survey and RP experts investigate unknown IR2 aperture bottleneck. Location from beam measurements: somewhere upstream of the TCTVB.4L2.

08:30 meeting

Wednesday 2nd November

  • Cryo recovery and access
  • 17:30 Cryo OK
  • 19:00 Injection
  • 23:00 Start of aperture measurement in Pt3
  • 03:00 beam dump due to losses when moving TCT.
  • No beam from SPS due to beam dump kicker problem

08:30 meeting & MD report

Cryogenics status

MD plan discussion

Tuesday 1st November

  • 01:00 Start of TDI and UFO MD
  • 09:00 Start UFO MD
  • 10:37 Lost cryogenics in point 8
  • Accesses during day and evening
  • TI8 stability tests overnight

08:30 meeting & MD report

Monday 31th October

  • 07:00 Start of Proton – Lead MD. First lead beam 2011. Ramped proton and lead beam. Re-phasing successful. Halloween night…

08:30 meeting
IR2 ions aperture (Jörg)